1.     What is the solubility product expression for silver carbonate?


a.      Ksp = [Ag⁺]² [CO₃²⁻]

b.     Ksp = [Ag⁺] [CO₃²⁻]²

c.      Ksp = 2 [Ag⁺] [CO₃²⁻]

d.     Ksp = ½ [Ag⁺] [CO₃²⁻]


 

      

3.     What is the molar solubility of the solution? (2 points)

a.      3.23 × 10⁻⁴

b.     2.24 × 10⁻³

c.      1.30 × 10⁻³

d.     1.28 × 10⁻⁴

 

    

5.     Calculate the mass of silver carbonate that will dissolve in 150.0 mL of water. (2 points)

a.      0.0134 g

b.     0.0927 g

c.      5.29 × 10⁻³ g

d.     0.0538 g

 


1
Expert's answer
2022-01-11T18:35:05-0500

Solution:

(1) What is the solubility product expression for silver carbonate?

silver carbonate = Ag2CO3

Ag2CO3(s) ⇌ 2Ag+(aq) + CO32−(aq)

Thus, the Ksp expression for Ag2CO3(s) is:

Ksp(Ag2CO3) = [Ag+][CO32−]

Answer (1):

a) Ksp = [Ag+][CO32−]


(3) What is the molar solubility of the solution?

Ksp(Ag2CO3) = 8.10×10−12 (the table value)

Ag2CO3(s) ⇌ 2Ag+(aq) + CO32−(aq)

___S__________2S_________S_______

Ksp(Ag2CO3) = [Ag+][CO32−] = [2S][S] = 4S3 = 8.10×10−12

4S3 = 8.10×10−12

S3 = 2.025×10−12

S = 0.0001265 = 1.27×10−4

The molar solubility of the solution is 1.27×10−4 M

Answer (3):

d) 1.28×10−4


(5) Calculate the mass of silver carbonate that will dissolve in 150.0 mL of water.

S(Ag2CO3) = 1.27×10−4 M = 1.27×10−4 mol/L

Volume of water = 150.0 mL = 0.15 L

The molar mass of Ag2CO3 is 275.7453 g/mol

Hence,

m(Ag2CO3) = S(Ag2CO3) × M(Ag2CO3) × V

m(Ag2CO3) = (1.27×10−4 mol L−1) × (275.7453 g mol−1) × (0.15 L) = 0.005253 g = 5.25×10−3 g

m(Ag2CO3) = 5.25×10−3 g

Answer (5):

c) 5.29×10−3 g